OpenStudy (anonymous):

let g(x) = 3x^2-5x+9 and assume that h ≠0 find g(2+h)-g(2)/h

OpenStudy (dape):

\[ g(2+h) = 3(2+h)^2-5(2+h)+9 \] and \[ g(2)/h = 11/h \] gives \[ g(2+h)-g(2)/h = 3(2+h)^2-5(2+h)+9 - 11/h \]

OpenStudy (accessdenied):

oh, you can do this, nvm. :x if you plug in the values from g(2+h) and g(2) and get it into a form where you can factor out an h from all terms, you can cancel it with the denominator to get the value